OpenStudy (anonymous):

Need help on solving this equation diamond equation 5 ? ? 10.5

OpenStudy (dumbcow):

what is this diamond equation you speak of ? :)

OpenStudy (anonymous):

5 ? ? 10.5 need the answwer to the question makrks. sorry delete buttom not working. question marks

OpenStudy (anonymous):

does any of this help? http://mathworld.wolfram.com/Diamond.html

OpenStudy (dumbcow):

i don't know what the question marks are referring to? i've never used "diamond equations"

OpenStudy (dumbcow):

@timo86m , thats prob little too advanced for what she is looking for ok so all that is given is 2 numbers and 2 unknowns and no function or relation?

OpenStudy (anonymous):

it could be like pascals triangle. It has no function but by popularity it is like the factors of a number or something.

OpenStudy (dumbcow):

could be ...@krazie32 any help?

OpenStudy (dumbcow):

help us help you :)

OpenStudy (anonymous):

I know the #s the in the middle have to equal 5 and anbd 10.5 by additiona and multipluication.

OpenStudy (anonymous):

maybe you mean pascals triangle

OpenStudy (dumbcow):

a+b = 10.5 ab = 5 a=10 b = 0.5

OpenStudy (anonymous):

or a+b=5 a*b=10.5